OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution?

OpenStudy (anonymous):

Asnaseer please help me afterrr

OpenStudy (asnaseer):

so you have 12L with 40% of it being alcohol - first step - have a party! but on a more serious note, 40% alcohol would be you have 0.4*12L of alcohol which comes to 4.8L. this means that amount of water in there is 12-4.8=7.2L. so, just to summarise, we start off with: 12L solution with 7.2L of water and 4.8L of alcohol now we need to add some amount of water to this - lets call that x. when we do that we will end up with: (12+x)L of solution, made up of: (7.2+x)L of water 4.8L of alcohol we want the amount of alcohol to represent 30% of the solution, so we can write: 4.8=0.3(12+x)=3.6+0.3x can you solve from here?

OpenStudy (anonymous):

4L ?

OpenStudy (asnaseer):

precisely - well done!

OpenStudy (anonymous):

Solve the equation: t/3 - 1/2 = 1

OpenStudy (anonymous):

help!!!!!